# ADT4Coupons: An Innovative Framework for Sequential Coupon Distribution in E-commerce

### v1

@@ 5.4 Inference Speed Analysis
Based on the results, ADT4Coupons framework achieves an inference speed suitable for industrial applications. In addition, we use 16 parallel models in the multi-user scenario, limited by CUDA memory. We anticipate that deployment in industrial environments, equipped with superior computing resources (e.g., larger GPU memory, parallel GPUs), will effectively mitigate this memory bottleneck.
6 Conclusion
In this paper, we propose ADT4Coupons framework, sequentially formulating coupon distribution strategies to enhance long-term revenue. Our unified framework combines general scenarios, sequential modeling and efficient iterative updates. Extensive experiments have been conducted to demonstrate the superiority of our framework.
Future directions includes integrating ADT4Coupons with externalities, such as incorporating user-to-user impact evaluation as input features. Also, we plan to encode coupons to prevent action space drift. This is in case we need to dynamically adapt coupon types (e.g., larger discounts) based on observed coupon performance.

### v2

@@ Introduction
Table 2: Model Inference Time
Based on the results, SACO framework achieves an inference speed suitable for industrial applications. In addition, we use 16 parallel models in the multi-user scenario, limited by CUDA memory. We anticipate that deployment in industrial environments, equipped with superior computing resources (e.g., larger GPU memory, parallel GPUs), will effectively mitigate this memory bottleneck.
Conclusion
In this paper, we propose SACO framework, sequentially formulating coupon distribution strategies to enhance long-term revenue. Our unified framework combines general scenarios, sequential modeling and efficient iterative updates. Extensive experiments have been conducted to demonstrate the superiority of our framework.
Future directions includes integrating SACO with externalities, such as incorporating user-to-user impact evaluation as input features. Also, we plan to encode coupons to prevent action space drift. This is in case we need to dynamically adapt coupon types (e.g., larger discounts) based on observed coupon performance.
Acknowledgments
This work was supported by National Natural Science Foundation of China (No.62472428), Public Computing Cloud, Renmin University of China, the fund for building world-class universities (disciplines) of Renmin University of China and ByteDance.